solve the following

(7)/(12ab) - (11)/(15a^2)
1
Expert's answer
2012-12-03T09:16:10-0500
(7)/(12ab) - (11)/(15a^2)=( 7*5a)/(12ab*5a)-(11*4b)/(15a^2*4b)=(35a-44b)/(60a^2b)